OpenStudy (anonymous):

Find all real zeros f(x)=3x(x-9)^2(x-6)^2

OpenStudy (phi):

they want you to find what x values make 3x(x-9)^2(x-6)^2=0

OpenStudy (phi):

and the key idea is if *any term* is 0, the product is 0

OpenStudy (phi):

ignoring the 3 which obviously can't be 0, you have 3 terms that might be 0, depending on what x is. can you solve for x for each term, when the term is set equal to zero ?

OpenStudy (anonymous):

Where do we start is my question

OpenStudy (phi):

does this part make sense if *any term* is 0, the product is 0 ?

OpenStudy (phi):

that is a yes or no question

OpenStudy (anonymous):

yes

OpenStudy (phi):

in 3x(x-9)^2(x-6)^2=0 list the 3 terms that could be zero: x (x-9)^2 (x-6)^2

OpenStudy (anonymous):

ok

OpenStudy (phi):

for each of those 3, we "solve" for x so that the term is 0 the first is easy , right ? the other two, by taking the square root of both sides in the equation \[ (x-9)^2 = 0 \]

OpenStudy (anonymous):

9 for that one

OpenStudy (phi):

yes, and because we can write (x-9)^2 as (x-9)(x-9)=0 we have two terms (both become 0 when x is 9) so point is, we call this a "repeated root"

OpenStudy (phi):

or sometimes we say, 9 with multiplicity 2 I assume you can find x for the other two terms?

OpenStudy (anonymous):

0,9,6 ?

OpenStudy (phi):

yes, and if we want to be picky, 0, 9 repeated twice, 6 repeated twice

OpenStudy (anonymous):

ok thank you!